OpenStudy (anonymous):

A satellite is in a 12,000 km radius Earth orbit (ME=5.98×10^24 kg). What is the speed of the satellite.

OpenStudy (anonymous):

I got it its GMm/r^2=mv^2/r v=sqrt(GM/r) (6.67e-11*5.98e24)/(6.37E6+12000E3) 4659 m/s